Hales TV Productions is considering producing a pilot for a comedy series in the hope of selling it to a major television ne(0) 5, -100 5, 50 5, 250 Agency Favorable 5, 150 5 5 150 5, 150 5. 5, -100 d. 50 250 5. No Agency Unfavorable 51 150 5, 150 S(iii) بهم -100 3 بم 1[ بنگر 0 Agency ار 10 d [ 13[(iv) 51 5. -100 6 50 250 Favorable نما 5 5, 150 d 150 S. 150 Agency 51 52 -100 4 8 50 s, 250 Unfavorable 4 5 5. 150 150 3 150b. What is the recommended decision if the agency opinion is not used? What is the expected value? Enter your answer in thousThe option for A is graph 1,2,3, or 4. For B C and D, the options are produce or sell. For F it is either yes or no.
Answer #1

(a) The correct decision tree is (iv)

0.06 S1 -100 -100 -100 di 0.25 S2 50 0 179 50 50 0.65 Favorable (F) 0.69 S3 250 0 179 250 250 d2 150 150 150 0.47 si Consult

(b)

For agency option not used: The recommended decision = Sell (d2)

Expected payoff = $ 150 thousand

(c)

EVPI = EVUPI - max. EMV = 0.2*Max(-100, 150) + 0.3*Max(50, 150) + 0.5*Max(250, 150) - 150 = $ 50 thousand

(d)

With agency information:

If favorable: Produce pilot (d1)
If unfavorable: Sell (d2)

(e)

EVSI = EV with the agency option - EV w/o the agency option = 168.85 - 150 = $ 18.85 thousand

(f)

Yes, becasue the maximum that can be paid is the EVSI i.e. $ 18.85 thousand